# Geonhwa Jeong

### v2

@@ 4.2 TASD-W: Applying TASD on weights
A better method to use TASD-W is using different TASD configurations for different layers, i.e. layer-wise TASD-W. The TASD transformations that can be covered by layer-wise TASD-W is a super-set of the TASD transformations in network-wise TASD-W. Usually, a pruned model with unstructured sparsity does not guarantee the same sparsity across layers, i.e. even though the overall sparsity is 95% for the model, different layers could have different sparsity degrees as shown in Figure 8. Unlike network-wise TASD-W where all the layers use the same TASD configuration, it is not straightforward how to choose a TASD configuration per layer as there could be numerous options per layer. To minimize the accuracy drop, it is crucial to reduce the number of dropped non-zeros after applying TASD, which would prefer conservative TASD configurations. On the other hand, to maximize the performance gain, it would be better to apply aggressive TASD configurations, which would be able to be translated into higher sparsity and efficiency gain.
To address this, we design and implement a greedy-based algorithm that optimizes across all layers. This greedy algorithm first measures the percentage of dropped non-zero elements of each TASD configuration for all layers and sorts the configuration-layer pairs by their percentage of dropped elements. Next, it greedily applies the TASD configuration based on the sorted order until the model quality is <99% of the original model (i.e., prioritize the option with the smallest dropped non-zeros). Since it only takes a single pass to all layers, the runtime overhead is trivial (a few seconds per model), while the extra training needed for structured sparse HW often takes hundreds of GPU hours [62].
@@ 4.3 TASD-A: Applying TASD on activations
To address this issue, we again leverage layer-wise TASD as it can tailor the TASD configuration to each layer. However, unlike the TASD-W, it is not feasible to test every option for each layer to find out the best options as the target tensors (activations) are dynamically generated.
Nonetheless, we find that a small set of calibration dataset (e.g., 1000 images for ImageNet [11]) can provide enough insights. As shown in Figure 8, while different layers have different sparsity degrees, for a particular layer, the activation sparsity degree remains in a small range across different input images. Therefore, TASDER takes calibration data as an input, so it can profile the given DNN model with the calibration data and collect the statistics (e.g., average, 99th percentile) about activation sparsity per layer.
To choose a TASD-A configuration for each layer, we use a sparsity-based selection method, instead of the non-zero-based method for TASD-W. We use a hyperparameter, $\alpha$, to tune the aggressiveness of the TASD approximation. For a given layer $L_{i}$ and the available configurations in the target HW (e.g., $H_{1},...,H_{4}$), we choose $C_{i}$ as $H_{j}$ where j is the largest integer where $S(L_{i})+\alpha>H_{j}$. If we use a larger $\alpha$, we choose the TASD configuration more aggressively (i.e. allowing more dropped non-zeros). We summarize the algorithm in 1.
Listing 1: The sparsity-degree-based TASD selection.
1target_sparsity = sparsity + alpha;
2if (target_sparsity > H[3])
3 return H[3];
4else if (target_sparsity > H[2])
5 return H[2];
6else if (target_sparsity > H[1])
7 return H[1];
8else if (target_sparsity > H[0])
9 return H[0];
Beyond sparsity: Supporting non-ReLU-based DNNs. ReLU-based DNNs naturally induce sparsity in activation tensors, so by collecting the sparsity statistics, TASD-A can find the appropriate configuration for each layer. However, for better accuracy, state-of-the-art DNNs have replaced ReLU with other activation functions, such as GeLU [21] and Swish [55], which do not induce any sparsity in activations making the activations dense. Thus, our sparsity-degree-based TASD selection (1) to choose TASD configuration for TASD-A for each layer would not work for those DNNs.
To address this, we investigate the distribution of the magnitude of all elements in the activation tensors from GeLU/Swish-based DNN. We found that, while no element in the tensor is exactly zero, a huge number of elements have tiny magnitude, compared to the range of magnitude for all elements. Therefore, we let TASD-A leverage this skewed distribution and collect the magnitude statistics. We introduce another heuristic, pseudo-density, which aims to preserve a fixed percentage (e.g., 99%) of the sum of all elements in a tensor, to determine the best TASD configuration for every layer. Using the pseudo-density for the non-ReLU-based DNNs, we can use the same sparsity-degree-based method (1) (i.e. by replacing sparsity to 1 - pseudo-density).
The approximating nature of TASD allows the system to also accelerate non-ReLU-based DNNs, while prior work that specifically targets activation sparsity cannot.
@@
Table 2: Supported sparse patterns with TTC-VEGETA.
@@ 4.4 Structured sparse HW for TASD
TASD works best when there are at least a few supported structured sparse patterns in the target sparse accelerator. While the TASDER optimizer is HW-agnostic, we propose to build on top of a recently proposed flexible structured sparse tensor accelerator to maximize the benefit. Inspired by previous structured sparse accelerators [36, 27, 46], we introduce TASD Tensor Core (TTC). We adopt a design similar to VEGETA [27] engine composed of multiple processing elements (PEs) while providing support for 1:8, 2:8, and 4:8 structured sparse patterns, and we call it TTC-VEGETA. With TASD and a limit of up to 2 terms, a TTC-VEGETA engine can support 7 out of all the N:8 patterns [^3] The 7:8 pattern needs 3 TASD terms and is rarely used in practice.as shown in Table 2 even though the original VEGETA supports only three sparse patterns.
Note that TTC can adopt other structured sparse designs, such as STC [72] with supports for 2:4 and dense, which we call TTC-STC. This would limit the flexibility in approximation using TASD compared to VEGETA-based TTC design, but TASDER is still able to optimize some layers. We explore the benefit of flexibility in Section 5.
Figure 11: TASD-HW composed of four TTCs.
@@
In Figure 11, we show the overall design of a TASD HW that is composed of four TTCs similar to the one used in the previous work [66]. The only modification we add on top of the N:M accelerator such as STC or VEGETA, is the TASD units (shown in the right part of Figure 11) that dynamically extract TASD terms from the activation tensor, similar to the DAP unit in the recent S2TA [37] accelerator. TASD-W can be applied offline through pre-processing since weights do not change during the runtime, but TASD-A requires the TASD unit as activations will be dynamically generated at runtime.
Given the computation latency on TTCs, the minimum number of TASD units per TTC required to hide the latency of TASD units depends on the mapping and TTC implementations. For example, each TTC-VEGETA with M=8 generates 16 (number of PE columns in each TTC) output elements per cycle (i.e. 2 blocks per cycle) as shown in Figure 12, which will be fed to TASD units. For an M-element block, a TASD unit sequentially extracts the largest values, so the decomposition takes up to M cycles for any TASD configuration as the sum of Ns in a TASD configuration cannot be larger than M.
The example in Figure 12 uses TASD configuration composed of 4:8 and 1:8, so it takes 5 cycles per block. At T1 (cycle 1), two blocks (Blk-1, Blk-2) will be produced from the PE array and Blk-1 and Blk-2 will be processed by TASD Unit 1 and TASD Unit 2, respectively (each cycle, two TASD units start execution). During T2-T5, Blk-1 and Blk-2 will be used to extract Decomposed Blocks, DBlk-1 and DBlk-2 for 4:8 Tiles. Then, during T6, DBlk-1 and DBlk-2 for 1:8 Tiles will be generated and stored. The decomposed blocks will be used as the inputs of the next layer. With 16 TASD units, a TTC-VEGETA can operate without stalls on the PE array due to the decomposition as a TASD unit is always guaranteed to be available after M cycles (i.e. by Little’s law, $16=2\times 8$). We also measure and present the area overhead for TASD units in Section 5.
@@
Decomposition-aware dataflow. In Figure 13, we show a mapping of a matrix multiplication with an approximated matrix A using a TASD configuration, 4:8 and 1:8 for the TTC. We first show how we tile the matrices and how they are mapped in the private register file and shared buffer of each TTC. When matrix $A$ is decomposed into two TASD terms, $A_{1}$, and $A_{2}$, the original matrix multiplication can be approximated as the sum of the two matrix multiplications and accumulation ($A_{1}\times B+A_{2}\times B$). As the two matrix operations share the same input $B$ and partial sum $C$, we keep B tiles in the L2 Scratchpad Memory (SMEM) and C tiles stationary in the L1 SMEM of TTC while changing decomposed A tiles to temporally reuse B and C tiles for data reuse (between timestep 1 and 2, timestep 3 and 4 in  Figure 13 (b)). For each accelerator tile, (i.e. for each timestep), we keep each element of A tile stationary in the register file of each PE for the temporal reuse, while the B and C elements are mapped correspondingly. By increasing the tile size for GEMM-N dimension, the reuse count for A tile at PE level could increase, which is limited by the size of the capacity of each SMEM. We swap C tiles at the very end to minimize the number of write-back operations to other levels. Although we maximize reuses for decomposed tiles, there is still unavoidable overhead such as reading C tiles again, but this is insignificant compared to the energy saving by skipping ineffectual computations using TASD. In Section 5.5, we quantify the energy overhead.
@@ 5.1 Methodology
HW Design HW Sparsity Support
@@
TTC-VEGETA-M4 1:4, 2:4 (TASD 1T) + 3:4 (TASD 2T)
TTC-VEGETA-M8 1:8, 2:8, 4:8 (TASD 1T) + 3:8, 5:8, 6:8 (TASD 2T)
TASD accelerates both sparse and dense DNNs without fine-tuning, so we evaluate TASD-W on sparse DNNs from SparseZoo [45] and TASD-A on dense DNNs from TorchVision [1]. We use a classic convolutional network, ResNet50 [20], and a transformer-based network, BERT [12], to illustrate detailed trade-offs between accuracy and performance. For the baseline HW, we compare against dense tensor core (TC) [46] and dual-side sparse tensor core (DSTC) [64] as representative dense and unstructured sparse accelerators. We configure these baselines as in the Sparseloop Artifacts [2]. We use 4 variants of TTC, based on STC and VEGETA, and with N:4 and N:8 designs, to show the extra benefits of TASD from the flexibility of the structured sparse hardware as summarized in Table 3. All designs use the same memory hierarchy and the same amount of PE (MACs) to ensure a fair comparison.
Model Weight Activation Layers Dimensions
Dense ResNet50 Dense Sparse L1: M784-N128-K1152
(ReLU-based) L2: M3136-N64-K576
Sparse ResNet50 Sparse Sparse L3: M196-K2304-N256
Dense BERT Dense Dense L1: M768-N128-K768
(GeLU-based) L2: M3072-N128-K768
@@
Table 4: Representative layers from the target workloads. L1, L2, and L3 are representative layers. We use "Overall" for the entire network.
@@
Refer to captionRefer to caption
Figure 16: Network-wise and Layer-wise TASD on ResNet50. Left: TASD-W. Right: TASD-A.
To evaluate the effectiveness of different TASD methods on a target accelerator, we develop TASDER as a framework to search for TASD transformations and calculate the accuracy of the model with each TASD transformation using PyTorch [51]. Following the requirement in MLPerf [56] inference benchmark, we only consider a model as valid with TASD if the model still achieves an accuracy higher than 99% of the accuracy from the original model. Next, we run each DNN layer with the given TASD series configuration using Sparseloop [66], a sparse tensor accelerator modeling framework to obtain per-layer results and aggregate the result for the entire network, which is consistent with prior accelerator simulation frameworks [49, 24, 31, 40, 58]. We simulate all layers in the networks, but to show per-layer results as well as the full network result ("Overall"), we also chose three representative layers (from early, mid, late) for each DNN as shown in Table 4. Also, to show the applicability of TASD for other sparse and dense DNNs, we evaluate the theoretical MACs reductions for another 8 DNNs, similar to prior work in pruning algorithms for structured sparse patterns [7, 62].
5.2 DNN acceleration with TASD
Figure 14 shows the energy-delay product (EDP) for the 4 workloads on various DNN accelerators, normalized to the dense TC.
Even though DSTC is able to exploit unstructured sparsity, the overhead of unstructured sparse acceleration (such as accessing accumulation buffer frequently) offsets the benefit and even outweighs the benefits when the workload has only one sparse operand or there is no sparse operand, causing 12% and 167% larger EDP for dense ResNet50 and dense BERT while reducing EDP by 55% for sparse BERT. DSTC works best for sparse ResNet50 and improves EDP by 87%, as both weight and activation tensors are unstructured sparse with a high sparsity degree (95% sparse weight). Overall, it is able to reduce 35% across all workloads on average.
Unlike DSTC, TASD-based TTC accelerators improve EDP over the TC baseline for all workloads. With the flexibility in sparsity patterns, TTC-VEGETA-M8 improves EDPs for all workloads, by 58%/61% for dense ResNet50/BERT and 83%/82% for sparse ResNet50/BERT. Even with only one fixed structured sparsity pattern, TTC-STC-M4 improves by 4%/32% for dense ResNet50/BERT and 49%/53% for sparse ResNet50/BERT. This result shows that TASD can effectively leverage structured sparse hardware for off-the-shelf dense and sparse DNNs with no fine-tuning, and the extra flexibility (increasing M) in the baseline accelerator increases the benefit.
Figure 15 provides more details in end-to-end latency and energy consumption for various designs. TTC-VEGETA-M8 is always the most energy-efficient design across all workloads and is slightly slower than DSTC only for sparse ResNet50 (by 22%). This result shows TASD provides a better overall tradeoff than unstructured sparse accelerators, especially considering their high area overheads.

### v3

@@ 4.2 TASD-W: Applying TASD on weights
A better method to use TASD-W is using different TASD configurations for different layers, i.e. layer-wise TASD-W. The set of TASD transformations that can be covered by layer-wise TASD-W is a super-set of the TASD transformations in network-wise TASD-W. Usually, a pruned model with unstructured sparsity does not guarantee the same sparsity across layers, i.e. even though the overall sparsity is 95% for the model, different layers could have different sparsity degrees as shown in Figure 6. Unlike network-wise TASD-W where all the layers use the same TASD configuration, it is not straightforward how to choose a TASD configuration per layer as there could be numerous options per layer. To minimize the accuracy drop, it is crucial to reduce the number of dropped non-zeros after applying TASD, which would prefer conservative TASD configurations. On the other hand, to maximize the performance gain, it would be better to apply aggressive TASD configurations, which would be able to be translated into higher sparsity and efficiency gain.
To address this, we design and implement a greedy-based algorithm that optimizes across all layers. This greedy algorithm first measures the percentage of dropped non-zero elements of each TASD configuration for all layers and sorts the configuration-layer pairs by their percentage of dropped elements. Next, it greedily applies the TASD configuration based on the sorted order until the model quality is $<$99% of the original model (i.e., prioritize the option with the smallest dropped non-zeros). Since it only takes a single pass to all layers, the runtime overhead is trivial (a few seconds per model), while the extra training needed for structured sparse HW often takes hundreds of GPU hours Sun et al. (2021).
@@ 4.3 TASD-A: Applying TASD on activations
To address this issue, we again leverage layer-wise TASD as it can tailor the TASD configuration to each layer. However, unlike the TASD-W, it is not feasible to test every option for each layer to find out the best options as the target tensors (activations) are dynamically generated. We find that a small set of calibration dataset (e.g., 1000 images for ImageNet Deng et al. (2009)) can provide enough insights. As shown in Figure 6, while different layers have different sparsity degrees, for a particular layer, the activation sparsity degree remains in a small range across different input images. Therefore, TASDER takes calibration data as an input, so it can profile the given DNN model with the calibration data and collect the statistics (e.g., average, 99th percentile) about activation sparsity per layer.
To choose a TASD-A configuration for each layer, we use a sparsity-based selection method, instead of the dropped-non-zero-based method for TASD-W. We use a hyperparameter, $\alpha$, to tune the aggressiveness of the TASD approximation. For a given layer $L_{i}$ and the available configurations in the target HW (e.g., $H_{1},...,H_{4}$), we choose $C_{i}$ as $H_{j}$ where $j$ is the largest integer where $S(L_{i})+\alpha>H_{j}$. If we use a larger $\alpha$, we choose the TASD configuration more aggressively (i.e. allowing more dropped non-zeros).
Beyond sparsity: Supporting non-ReLU-based DNNs. For better accuracy, state-of-the-art DNNs have replaced ReLU with other activation functions, such as GeLU Hendrycks & Gimpel (2016) and Swish Ramachandran et al. (2017), which do not induce any sparsity in activations making the activations dense. Thus, our sparsity-degree-based TASD selection to choose TASD configuration for TASD-A for each layer would not work for those DNNs.
Table 2: Supported sparse patterns with TTC-VEGETA.
@@
To address this, we investigate the distribution of the magnitude of all elements in the activation tensors from GeLU/Swish-based DNN. We found that, while no element in the tensor is exactly zero, a huge number of elements have tiny magnitude, compared to the range of magnitude for all elements. Therefore, we let TASD-A leverage this skewed distribution and collect the magnitude statistics. We introduce another heuristic, pseudo-density, which aims to preserve a fixed percentage (e.g., 99%) of the sum of all elements in a tensor, to determine the best TASD configuration for every layer. Using the pseudo-density for the non-ReLU-based DNNs, we can use the same sparsity-degree-based method (i.e. by replacing sparsity to 1 - pseudo-density). The approximating nature of TASD allows the system to also accelerate non-ReLU-based DNNs, while prior work that specifically targets activation sparsity cannot.
Figure 9: TASD-HW composed of four TTCs.
@@ 4.4 Structured sparse HW for TASD
TASD works best when there are at least a few supported structured sparse patterns in the target sparse accelerator. While the TASDER optimizer is HW-agnostic, we propose to build on top of a recently proposed flexible structured sparse tensor accelerator to maximize the benefit. Inspired by previous structured sparse accelerators NVIDIA (2020a); Liu et al. (2021); Jeong et al. (2023), we introduce TASD Tensor Core (TTC). We adopt a design similar to VEGETA Jeong et al. (2023) engine composed of multiple processing elements (PEs) while providing support for 1:8, 2:8, and 4:8 structured sparse patterns, and we call it TTC-VEGETA. With TASD and a limit of up to 2 terms, a TTC-VEGETA engine can support 7 out of all the N:8 patterns as shown in Table 2 even though the original VEGETA supports only three sparse patterns. Note that TTC can adopt other structured sparse designs, such as STC Zhu et al. (2019) with supports for 2:4 and dense, which we call TTC-STC. This would limit the flexibility in approximation using TASD compared to VEGETA-based TTC design, but TASDER is still able to optimize some layers. We explore the benefit of flexibility in section 5.
@@
In Figure 9, we show the overall design of a TASD HW composed of four TTCs similar to the one used in the previous work Wu et al. (2022b). The modification we add on top of the N:M accelerator, such as STC or VEGETA, is the TASD unit (shown in the right part of Figure 9) that dynamically extracts TASD terms from the activation tensor, similar to the DAP unit in S2TA Liu et al. (2022b). TASD-W can be applied offline through pre-processing since weights do not change during runtime, but TASD-A requires the TASD unit as activations will be dynamically generated at runtime.
Given the computation latency on TTCs, the minimum number of TASD units per TTC to hide the latency of TASD units depends on the mapping and TTC implementations. For example, each TTC-VEGETA with M=8 generates 16 (number of PE columns in each TTC) output elements per cycle (i.e. 2 blocks per cycle) as shown in Figure 10, which will be fed to TASD units. For an M-element block, a TASD unit sequentially extracts the largest values, so the decomposition takes up to M cycles as the sum of Ns in a TASD configuration cannot be larger than M.
The example in Figure 10 uses TASD configuration composed of 4:8 and 1:8, so it takes 5 cycles per block. At T1 (cycle 1), two blocks (Blk-1, Blk-2) will be produced from the PE array, and Blk-1 and Blk-2 will be processed by TASD Unit 1 and TASD Unit 2, respectively (each cycle, two TASD units start execution). During T2-T5, Blk-1 and Blk-2 will be used to extract Decomposed Blocks, DBlk-1 and DBlk-2 for 4:8 Tiles. Then, during T6, DBlk-1 and DBlk-2 for 1:8 Tiles will be generated and stored. The decomposed blocks will be used as the inputs of the next layer. With 16 TASD units, a TTC-VEGETA can operate without stalls on the PE array due to the decomposition as a TASD unit is always guaranteed to be available after M cycles (i.e. by Little’s law, $16=2\times 8$). We present the area overhead for TASD units in section 5.
@@
Decomposition-aware dataflow. In Figure 11, we show a mapping of a matrix multiplication with an approximated matrix A using a TASD configuration, 4:8 and 1:8 for the TTC. We first show how we tile the matrices and how they are mapped in the private register file and shared buffer of each TTC. When matrix $A$ is decomposed into two TASD terms, $A_{1}$, and $A_{2}$, the original matrix multiplication can be approximated as the sum of the two matrix multiplications and accumulation ($A_{1}\times B+A_{2}\times B$). As the two matrix operations share the same input $B$ and partial sum $C$, we keep B tiles in the L2 Scratchpad Memory (SMEM) and C tiles stationary in the L1 SMEM of TTC while changing decomposed A tiles to temporally reuse B and C tiles for data reuse (between timestep 1 and 2, timestep 3 and 4 in  Figure 11 (b)). For each accelerator tile, (i.e. for each timestep), we keep each element of A tile stationary in the register file of each PE for the temporal reuse, while the B and C elements are mapped correspondingly. By increasing the tile size for GEMM-N dimension, the reuse count for A tile at PE level could increase, which is limited by the size of the capacity of each SMEM. We swap C tiles at the very end to minimize the number of write-back operations to other levels. Although we maximize reuses for decomposed tiles, there is still unavoidable overhead such as reading C tiles again, but this is insignificant compared to the energy saving by skipping ineffectual computations using TASD. In subsection 5.4, we quantify the energy overhead.
@@ 5.1 Methodology
TASD accelerates both sparse and dense DNNs without fine-tuning, so we evaluate TASD-W on sparse DNNs from SparseZoo Neuralmagic (2023) and TASD-A on dense DNNs from TorchVision PyTorch (2023). We use a classic convolutional network, ResNet50 He et al. (2016), and a transformer-based network, BERT Devlin et al. (2019). For the baseline HW, we compare against dense tensor core (TC) NVIDIA (2020a) and dual-side sparse tensor core (DSTC) Wang et al. (2021) as representative dense and unstructured sparse accelerators. We configure these baselines as in the Sparseloop Artifacts Wu et al. (2022a). We use 4 variants of TTC, based on STC and VEGETA with N:4 and N:8 designs, to show the extra benefits of TASD from the flexibility of the structured sparse hardware as summarized in Table 3. All designs use the same memory hierarchy and the same amount of PEs (MACs) to ensure a fair comparison. We clarify that we do not exploit both sparsities concurrently for skipping computations. We either use TASD-W or TASD-A depending on the workloads since supporting both sparsities requires non-trivial area/energy costs Wu et al. (2021; 2023) and is not compatible to approximate with TASD.
We develop TASDER as a framework to search for TASD transformations and calculate the accuracy of the model with each TASD transformation using PyTorch Paszke et al. (2019). Following the requirement in MLPerf Reddi et al. (2020) inference benchmark, we only consider a model as valid with TASD if the model achieves an accuracy higher than 99% of the accuracy of the original model. Next, we run each DNN layer with the given TASD series configuration using Sparseloop Wu et al. (2022b), a sparse accelerator modeling framework to obtain per-layer results and aggregate the results for the entire network, which is consistent with prior accelerator simulation frameworks Parashar et al. (2019); Huang et al. (2021); Kwon et al. (2020); Mei et al. (2021); Samajdar et al. (2020). Sparseloop estimates the performance and energy consumption of the given HW by analyzing data movements across different memory hierarchy levels and processing elements, and actual computation based on the given dataflow. We simulate all layers in the networks (marked as “Overall”), but to show per-layer results additionally, we choose three representative layers (from early, mid, late) as shown in Table 4.
HW Design Sparsity Support from HW
@@
TTC-VEGETA-M4
1:4, 2:4 (TASD 1T)
+ 3:4 (TASD 2T)
TTC-VEGETA-M8
1:8, 2:8, 4:8 (TASD 1T)
+ 3:8, 5:8, 6:8 (TASD 2T)
Table 4: Representative layers from the target workloads. L1, L2, and L3 are representative layers. RN and Act. refer to ResNet and Activation, respectively.
Model Weight Act. Layers Dimensions
Dense RN50
(ReLU-based)
Dense Sparse
L1: M784-N128-K1152
L2: M3136-N64-K576
Sparse RN50 Sparse Sparse L3: M196-K2304-N256
Dense BERT
(GeLU-based)
Dense Dense
L1: M768-N128-K768
L2: M3072-N128-K768
@@
We use energy-delay product (EDP), latency, and energy for the metrics. We also provide additional experiment results on the theoretical MACs reductions for other various DNNs, comparison against structured sparse accelerators, and TASD on a real system.
5.2 DNN acceleration with TASD
Figure 12 shows the EDP for the 4 workloads on various DNN accelerators, normalized to the dense TC. Even though DSTC is able to exploit unstructured sparsity, the overhead of unstructured sparse acceleration (such as accessing accumulation buffer frequently) offsets the benefit and even outweighs the benefits when the workload has only one sparse operand or there is no sparse operand, causing 12% and 167% larger EDP for dense ResNet50 and dense BERT while reducing EDP by 55% for sparse BERT. DSTC works best for sparse ResNet50 and improves EDP by 87%, as both weight and activation tensors are unstructured sparse with a high sparsity degree (95% sparse weight). Overall, it is able to reduce EDP by 35% across all workloads on average.
Unlike DSTC, TASD-based TTC accelerators improve EDP over the TC baseline for all workloads. With the flexibility in sparsity patterns, TTC-VEGETA-M8 improves EDP for all workloads, by 58%/61% for dense ResNet50/BERT and 83%/82% for sparse ResNet50/BERT. Even with a single fixed pattern, TTC-STC-M4 improves by 4%/32% for dense ResNet50/BERT and 49%/53% for sparse ResNet50/BERT. This result shows that TASD can effectively leverage structured sparse hardware for off-the-shelf dense and sparse DNNs with no fine-tuning, and the extra flexibility (increasing M) in the baseline accelerator increases the benefit.
Figure 13 provides end-to-end latency and energy consumption for various designs. TTC-VEGETA-M8 is the most energy-efficient across all workloads and is slightly slower than DSTC only for sparse ResNet50. This result shows TASD provides a better overall tradeoff than unstructured sparse accelerators, considering their high area overheads.
@@ 5.2 DNN acceleration with TASD
Figure 14: Network-wise and Layer-wise TASD on ResNet50. Upper: TASD-W. Lower: TASD-A.